# Spectral Signature Mapping from RGB Imagery for Terrain-Aware Navigation

### v1

@@ Abstract
Successful navigation in outdoor environments requires accurate prediction of the physical interactions between the robot and the terrain. To this end, several methods rely on geometric or semantic labels to classify traversable surfaces. However, such labels cannot distinguish visually similar surfaces that differ in material properties. Spectral sensors enable inference of material composition from surface reflectance measured across multiple wavelength bands. Although spectral sensing is gaining traction in robotics, widespread deployment remains constrained by the need for custom hardware integration, high sensor costs, and compute-intensive processing pipelines. In this paper, we present RGB Image to Spectral Signature Neural Network (RS-Net), a deep neural network designed to bridge the gap between the accessibility of RGB sensing and the rich material information provided by spectral data. RS-Net predicts spectral signatures from RGB patches, which we map to terrain labels and friction coefficients. The resulting terrain classifications are integrated into a sampling-based motion planner for a wheeled robot operating in outdoor environments. Likewise, the friction estimates are incorporated into a contact-force–based MPC for a quadruped robot navigating slippery surfaces. Thus, we introduce a framework that learns the task-relevant physical property once during training and thereafter relies solely on RGB sensing at test time. The code is available at https://github.com/prajapatisarvesh/RS-Net.
@@ I Introduction
Autonomous robots are increasingly deployed in everyday settings, ranging from self-driving taxis [1] and search-and-rescue missions [2, 3] to wildfire prevention [4, 5]. In such unstructured environments, reliable autonomy demands more than obstacle avoidance. It requires precise reasoning about how terrain properties influence motion. For example, vehicles must modulate braking on icy roads, and off-road platforms should bypass dense swamps to avoid entrapment. These scenarios show that perception must move beyond geometry and semantics toward reliable estimates of robot–terrain interactions.
Several off-road motion planning pipelines use RGB cameras to identify terrain from images [6, 7, 8]. In some cases, visually similar surfaces with different physical properties, such as ice on asphalt, may be mislabeled, leading to invalid traversability cost maps. Depth cameras and LiDAR are often used to estimate the ease of motion over a surface [9, 10]. However, the robot must first drive the terrain to create a dataset, which risks hardware damage and necessitates tuning specific to the operating site.
@@
RGB cameras are inexpensive, widely available, and already standard in robotic perception pipelines such as object detection [14] and tracking [15]. Compared to hyperspectral systems, they are cheaper, lighter, and more power efficient, which simplifies integration on mobile platforms. Advances in deep learning [16] now allow RGB imagery to approximate measurements traditionally obtained from more information-dense sensors.
Motivated by this, we seek to retain the deployment advantages of RGB cameras while recovering spectral sensor features. We introduce RGB Image to Spectral Signature Neural Network or RS-Net, a deep neural network architecture trained on spectral data collected from diverse materials. It maps RGB image patches to their corresponding spectral signatures. These estimates are passed to a lightweight feedforward neural network whose weights are fine-tuned for the target physical property. We retrain this network once per task, enabling the same neural network architecture to perform terrain classification and friction estimation. Our entire inference pipeline runs at approximately 5 Hz, making it suitable for real-time robotic applications. Fig. LABEL:fig:paper_intro outlines the proposed architecture.
@@
II Related Work
Inaccurate estimation of the coefficient of friction (CoF) can lead to slip and skid in mobile robots, resulting in degraded motion planning performance [17]. Many methods that rely on RGB cameras estimate friction via terrain classification. These approaches depend heavily on the quality of segmentation masks and are vulnerable to visual aliasing and mislabeling [6, 8, 18]. Moreover, models trained in one region, season, or camera setup may encounter significantly different illumination, exposure, or texture statistics at deployment, leading to poor generalization. Recent retrieval-augmented vision–language models avoid the need for paired image–friction labels by using language supervision for CoF estimation [19]. However, this comes at the cost of increased inference time. To address these limitations, we bypass semantic segmentation entirely and infer the friction coefficient from predicted spectral signatures. Our approach employs lightweight convolutional and fully connected networks, enabling inference at approximately 5 Hz. This supports real-time deployment while avoiding the computational overhead associated with retrieval-based language models.
@@ II Related Work
Spectral and hyperspectral sensors capture material-specific reflectance by measuring how surfaces respond across a broad range of wavelengths [26]. This enables fine-grained discrimination between materials that appear visually similar [12, 27, 28]. However, real-world deployment faces several challenges [3]. These include high sensor cost, complex mechanical integration on mobile robots, and large memory requirements due to the high-dimensional nature of the underlying data. We address the limitations of spectral sensing by predicting spectral profiles directly from RGB terrain patch. This removes the need for expensive, high-bandwidth sensors while leveraging compact, low-cost RGB cameras already integrated into most robotic platforms. The reduced data dimensionality enables efficient storage and real-time inference, making the approach practical for large-scale, field-deployable systems.
Recent traversability estimation methods are beginning to move beyond pure geometry by incorporating richer terrain cues. TERP [29] constructs stability-based costs from elevation maps, while TRIP [10] predicts traversability from steppability analysis on dense 3D reconstructions. However, both these methods lack explicit modeling of material properties and can still misclassify low-friction surfaces. EVORA [9] learns traction distributions from exteroceptive sensing and incorporates these estimates into a CVaR-aware motion planner. However, it relies on supervised learning from physical interaction data collected over rough terrain, which limits its scalability. In contrast, we propose a unified approach that estimates planner-agnostic material properties by fine-tuning a lightweight fully connected head on top of spectral inference network. This allows the same backbone to be reused across motion planning tasks with minimal retraining.
@@ III-A RS-Net Architecture and Task-Specific Prediction Head
Figure 2: Overview of our motion planning pipeline. SAM extracts image patches from the RGB input, which are passed through RS-Net followed by a Fully Connected Network to predict per-patch friction estimates or terrain labels. These predictions inform friction constraints or terrain-aware cost maps for downstream motion planning algorithms.
@@ III-B Neural Network Training Workflow
We train on the VAST dataset released by Hanson et al. [28], which contains high-resolution RGB images of terrain surfaces, paired spectral reflectance profiles, and 9-DoF IMU data across 11 terrain types under varied lighting conditions. RS-Net is trained on six terrain classes: asphalt, brick, grass, ice, sand, and tile. It is also evaluated on five held-out classes: carpet, concrete, gravel, mulch, and turf, to assess generalization.
@@ IV Experiments and Results
We perform three sets of experiments to evaluate RS-Net. First, we assess its ability to reconstruct spectral signatures from RGB images through standalone spectral estimation. Next, we integrate the friction prediction module into a quadruped robot’s control stack and evaluate its impact on navigation performance. Finally, we deploy RS-Net’s terrain classification outputs on a wheeled robot to guide outdoor navigation.
@@ IV-A Evaluation of RGB-to-Spectral Mapping
A key motivation for our approach is to address the failure of appearance based terrain classification methods in distinguishing between materials that look similar but differ in physical properties. Fig. 3 illustrates this using two terrain patches, asphalt and gravel, that appear nearly identical in RGB imagery. Despite this visual similarity, RS-Net reconstructs distinct spectral signatures for each. Asphalt exhibits consistently higher photon counts than gravel, allowing clear separation in the spectral domain. This demonstrates that RS-Net learns to extract fine-grained visual cues beyond texture to infer material composition. This capability is essential for downstream planning and control tasks that require reliable reasoning about terrain properties in visually ambiguous environments.
Figure 4: Representative surface textures used in simulation. From left to right: asphalt, mud, and ice.
@@ IV-B Quadrupedal Robot Navigation
where $f_{x}$, $f_{y}$, and $f_{z}$ are the contact forces along the respective axes. The shape of this cone, and therefore the feasible set of forces, is directly defined by $\mu$ [34], making accurate friction estimation critical for stable locomotion. The Linear Model Predictive Controller (LMPC) we use is based on [17]. The MPC parameters used for simulation and hardware experiments are shown in Table I. The controller tracks a desired center of mass height of 0.32 meters and a forward velocity of 0.5 $m/s$, while enforcing the friction cone constraint during optimization. We evaluate our method in both simulation and hardware. In each case, RGB images are segmented into patches using SAM. A friction estimate is computed for each patch, and the minimum predicted value is used to inform the LMPC.
TABLE I: LMPC parameters used for quadruped navigation.
@@ IV-B1 Monte Carlo Evaluation in Simulation
TABLE II: Results from 1000 Monte-Carlo Simulations.
@@
As shown in Table II, for the trotting gait, our framework performs better than simply using a fixed but small friction value. This improvement comes from the controller’s ability to adapt the friction cone based on the predicted terrain properties, rather than applying the same force limits across all surfaces. To further evaluate stability under low-friction conditions, we test the controller on an icy surface while tracking the center of mass height at 0.32 meters, as shown in Fig. 5. With active friction estimation via RS-Net, the robot maintains stability throughout the trial, whereas the fixed-friction baseline loses balance and falls within 3 seconds.
@@ IV-C Wheeled Robot Navigation
We compare two planning conditions: (i) Baseline MPPI, which uses geometry-only costs and ignores material cues, and (ii) RS-Net + MPPI, which integrates a terrain-aware cost layer based on RS-Net prediction. In repeated trials, both planners use identical MPPI hyperparameters. The baseline selects the shortest diagonal path through grass, minimizing time-to-goal but traversing less reliable terrain. In contrast, as shown in Fig. 8, the RS-Net-enabled planner avoids grass entirely, favoring longer asphalt-only routes. This yields comparable time-to-goal while significantly reducing the risk of getting stuck.
@@ V Limitations
Fig. 9 reports the spectral signature predicted by RS-Net for turf terrain, which was not included in the training set. The overall spectral shape is preserved, with peaks and dips occurring at similar wavelengths, but the predicted magnitude is consistently biased. This indicates a limitation of the current model. Although the trend is captured, the prediction is not precisely calibrated for unseen classes. To address this, we aim to explore alternate feature extractors beyond DenseNet [38] and perform additional tuning on larger and more diverse spectral datasets [39]. Additionally, we aim to explore pixel-wise spectral inference [40] in place of the current patch based approach. This may lead to more fine grained predictions and improved spatial resolution of terrain estimates.
@@ VI Conclusion
We present a modular architecture for estimating terrain properties to support motion planning in mobile robots. This is achieved by training a neural network to predict spectral signatures directly from RGB images, enabling low cost sensing without the need for dedicated spectral hardware. A simple training mechanism allows the output to be adapted for different physical properties, such as terrain class and friction coefficient. In future work, we plan to extend this framework beyond navigation to tasks such as manipulation, and to infer additional physical properties relevant to robot–environment interaction, including moisture content [41] and terrain compliance [42].

### v2

@@ Abstract
Successful navigation in outdoor environments requires accurate prediction of the physical interactions between the robot and the terrain. Many prior methods rely on geometric or semantic labels to classify traversable surfaces. However, such labels cannot distinguish visually similar surfaces that differ in material properties. Spectral sensors enable inference of material composition from surface reflectance measured across multiple wavelength bands. Although spectral sensing is gaining traction in robotics, widespread deployment remains constrained by the need for custom hardware integration, high sensor costs, and compute-intensive processing pipelines. In this paper, we present the RGB Image to Spectral Signature Neural Network (RS-Net), a deep neural network designed to bridge the gap between the accessibility of RGB sensing and the rich material information provided by spectral data. RS-Net predicts spectral signatures from RGB patches, which we map to terrain labels and friction coefficients. The resulting terrain classifications are integrated into a sampling-based motion planner for a wheeled robot operating in outdoor environments. Likewise, the friction estimates are incorporated into a contact-force–based MPC for a quadruped robot navigating slippery surfaces. Overall, our framework learns the task-relevant physical properties offline during training and thereafter relies solely on RGB sensing at run time. The code is available at https://github.com/prajapatisarvesh/RS-Net.
@@ I Introduction
Robots are increasingly deployed in everyday settings, ranging from self-driving cars [1] and search-and-rescue missions [2] to wildfire prevention [3, 4]. In such unstructured environments, reliable autonomy demands more than obstacle avoidance. It requires precise reasoning about how terrain properties influence motion. For example, vehicles must modulate braking on icy roads, and off-road platforms should bypass dense swamps to avoid entrapment. These scenarios show that perception must move beyond geometry and semantics toward reliable estimates of robot–terrain interactions.
Several off-road motion planning pipelines use RGB cameras to identify terrain from images [5, 6]. In some cases, visually similar surfaces with different physical properties, such as ice on asphalt, may be mislabeled, leading to invalid traversability cost maps. Depth cameras and LiDAR are often used to estimate the ease of motion over a surface [7, 8]. However, the robot must first drive the terrain to create a dataset, which risks hardware damage and necessitates tuning specific to the operating site.
@@
RGB cameras are inexpensive, widely available, and already standard in robotic perception pipelines such as object detection [12] and tracking [13]. Compared to hyperspectral systems, they are cheaper, lighter, and more power efficient, which simplifies integration on mobile platforms. Advances in deep learning now allow RGB imagery to approximate measurements traditionally obtained from more information-dense sensors.
Figure 1: Our approach takes RGB image patches (Left) as input and predicts their spectral signatures (Middle). These predicted spectra are then mapped to terrain classes by tuning the final fully connected layer (Right). The resulting terrain predictions inform navigation. The robot selects a path along the asphalt road, which is easier to traverse than the surrounding dense grass.
Motivated by this, we seek to retain the deployment advantages of RGB cameras while recovering spectral sensor features. We introduce the RGB Image to Spectral Signature Neural Network or RS-Net, a deep neural network architecture trained on spectral data collected from diverse materials. It maps RGB image patches to their corresponding spectral signatures. These estimates are passed to a lightweight feedforward neural network whose weights are fine-tuned for the target physical property. We retrain this network once per task, enabling the same neural network architecture to perform terrain classification and friction estimation. Our entire inference pipeline runs at approximately 5 Hz, making it suitable for real-time robotic applications. Fig. 1 outlines the proposed architecture.
@@
II Related Work
Inaccurate estimation of the coefficient of friction (CoF) can lead to slip and skid in mobile robots, resulting in degraded motion planning performance [14, 15]. Many methods that rely on RGB cameras estimate friction via terrain classification. These approaches depend heavily on the quality of segmentation masks and are vulnerable to visual aliasing and mislabeling [5]. Moreover, models trained in one region, season, or camera setup may encounter significantly different illumination, exposure, or texture statistics at deployment, leading to poor generalization. Recent retrieval-augmented vision–language models avoid the need for paired image–friction labels by using language supervision for CoF estimation [16]. However, this comes at the cost of increased inference time. To address these limitations, we bypass semantic segmentation entirely and infer the friction coefficient from predicted spectral signatures. Our approach employs lightweight convolutional and fully connected networks, enabling inference at approximately 5 Hz. This supports real-time deployment while avoiding the computational overhead associated with retrieval-based language models.
@@ II Related Work
Spectral and hyperspectral sensors capture material-specific reflectance by measuring how surfaces respond across a broad range of wavelengths [23]. This enables fine-grained discrimination between materials that appear visually similar [10, 24]. However, real-world deployment faces several challenges [2]. These include high sensor cost, complex mechanical integration on mobile robots, and large memory requirements due to the high-dimensional nature of the underlying data. We address the limitations of spectral sensing by predicting spectral profiles directly from RGB terrain patch. This removes the need for expensive, high-bandwidth sensors while leveraging compact, low-cost RGB cameras already integrated into most robotic platforms. The reduced data dimensionality enables efficient storage and real-time inference, making the approach practical for large-scale, field-deployable systems.
Recent traversability estimation methods are beginning to move beyond pure geometry by incorporating richer terrain cues. TERP [25] constructs stability-based costs from elevation maps, while TRIP [8] predicts traversability from steppability analysis on dense 3D reconstructions. However, both these methods lack explicit modeling of material properties and can still misclassify low-friction surfaces. EVORA [7] learns traction distributions from exteroceptive sensing and incorporates these estimates into a Conditional Value at Risk (CVaR)-aware motion planner. However, it relies on supervised learning from physical interaction data collected over rough terrain, which limits its scalability. In contrast, we propose a unified approach that estimates planner-agnostic material properties by fine-tuning a lightweight fully connected head on top of spectral inference network. This allows the same backbone to be reused across motion planning tasks with minimal retraining.
@@ III-A RS-Net Architecture and Task-Specific Prediction Head
Figure 3: Overview of our motion planning pipeline. Segment Anything (SAM) extracts image patches from the RGB input, which are passed through RS-Net followed by a Fully Connected Network to predict per-patch friction estimates or terrain labels. These predictions inform friction constraints or terrain-aware cost maps for downstream motion planning algorithms.
@@ III-B Neural Network Training Workflow
We train on the VAST dataset [24], which contains high-resolution RGB images of terrain surfaces, paired spectral reflectance profiles, and 9-DoF IMU data across 11 terrain types under varied lighting conditions. RS-Net is trained on six terrain classes: asphalt, brick, grass, ice, sand, and tile. It is also evaluated on five held-out classes: carpet, concrete, gravel, mulch, and turf, to assess generalization.
@@ IV Experiments and Results
We perform three sets of experiments to evaluate RS-Net. First, we assess its ability to reconstruct spectral signatures from RGB images. Next, we integrate the friction prediction module into a quadruped robot’s control stack and evaluate its impact on navigation performance. Finally, we deploy RS-Net’s terrain classification outputs on a wheeled robot to guide outdoor navigation.
@@ IV-A Evaluation of RGB-to-Spectral Mapping
TABLE I: MSE for Spectral Profile Prediction.
Model Score
SRCNN $\pm0.01$
[…]
A key motivation for our approach is to address the failure of appearance based terrain classification methods in distinguishing between materials that look similar but differ in physical properties. Fig. 4 illustrates this using two terrain patches, asphalt and gravel, that appear nearly identical in RGB imagery. Despite this visual similarity, RS-Net reconstructs distinct spectral signatures for each. Asphalt exhibits consistently higher photon counts than gravel, allowing clear separation in the spectral domain. This demonstrates that RS-Net learns to extract fine-grained visual cues beyond texture to infer material composition. This capability is essential for downstream motion planning tasks that require reliable reasoning about terrain properties in visually ambiguous environments.
To compare spectral signature prediction performance, we used 30% of the VAST dataset as the validation set. We evaluated our method against two popular networks: SRCNN [30] and U-Net [31], where we removed their final layers and added fully connected heads to perform spectral regression. As seen in Table I RS-Net achieves lower MSE across all terrain classes compared to both baselines.
Figure 5: Representative surface textures used in simulation. From left to right: asphalt, brick, and ice.
@@ IV-B Quadrupedal Robot Navigation
where $f_{x}$, $f_{y}$, and $f_{z}$ are the contact forces along the respective axes. The shape of this cone, and therefore the feasible set of forces, is directly defined by $\mu$ [32], making friction estimation critical for stable locomotion. The Linear Model Predictive Controller (LMPC) we use is based on [14]. The MPC parameters used for simulation and hardware experiments are shown in Table II. The controller tracks a desired center of mass height of 0.32 m and a forward velocity of 0.5 $m/s$, while enforcing the friction cone constraint during optimization. We evaluate our method in both simulation and hardware. In each case, RGB images are segmented into patches using SAM. A friction estimate is computed for each patch, and the minimum predicted value is used to inform the LMPC.
TABLE II: LMPC parameters used for quadruped navigation.
@@ IV-B1 Monte Carlo Evaluation in Simulation
TABLE III: Results from 1000 Monte-Carlo Simulations.
@@
As shown in Table III, for the trotting gait, our framework performs better than simply using a fixed but small friction value. This improvement comes from the controller’s ability to adapt the friction cone based on the predicted terrain properties, rather than applying the same force limits across all surfaces. To further evaluate stability under low-friction conditions, we test the controller on an icy surface while tracking the center of mass height at 0.32 meters, as shown in Fig. 6. With active friction estimation via RS-Net, the robot maintains stability throughout the trial, whereas the fixed-friction baseline loses balance and falls within 3 seconds.
@@ IV-C Wheeled Robot Navigation
To assess how well RS-Net performs terrain classification, we evaluate it on the VAST validation set and visualize the results in Fig. 8. While several terrain classes such as ice, grass, and tile are predicted with high accuracy, others like sand and brick show moderate confusion due to similarly shaped but spatially offset spectral profiles [34].
@@
We compare two planning conditions: (i) Baseline MPPI, which uses geometry-only costs and ignores material cues, and (ii) RS-Net + MPPI, which integrates a terrain-aware cost layer based on RS-Net prediction. Both planners use identical MPPI hyperparameters. The baseline selects the shortest diagonal path through grass, minimizing time-to-goal but traversing less reliable terrain. In contrast, as shown in Fig. 10, the RS-Net-enabled planner avoids grass entirely, favoring longer asphalt-only routes. This yields comparable time-to-goal while significantly reducing the risk of getting stuck.
@@ V Limitations and Future Work
Fig. 11 reports the spectral signature predicted by RS-Net for turf terrain, which was not included in the training set. The overall spectral shape is preserved, but the predicted magnitude is consistently biased. To mitigate this, we plan to investigate vision encoders with richer feature representations, such as Vision Transformer [37], and train on larger and more diverse datasets [38].
Instead of relying on a single patch, prediction accuracy may benefit from fusing information across multiple patches or viewpoints. This approach can help the model better capture the bidirectional reflectance distribution function (BRDF), which encodes how material properties vary with view and illumination. By incorporating such fine-grained physical cues, the model may reduce the discrepancy between predicted and ground truth spectra.
@@ VI Conclusion
We present a modular architecture for estimating terrain properties to support motion planning in mobile robots. This is achieved by training a neural network to predict spectral signatures directly from RGB images, enabling low cost sensing without the need for dedicated spectral hardware. A simple training mechanism allows the output to be adapted for different physical properties, such as terrain class and friction coefficient. In future work, we plan to extend this framework beyond navigation to tasks such as manipulation, and to infer additional physical properties relevant to robot–environment interaction, including moisture content and terrain compliance [39].